OpenStudy (anonymous):

Solve by factoring: x^2+5x+6=0

OpenStudy (anonymous):

@seesawn @wolf1728

OpenStudy (anonymous):

(x+3)(x+2

OpenStudy (wolf1728):

x^2+5x+6=0 What I do when factoring. I look at the third term and see what factors can comprise it. Hmmmmmmm seems as though Griff has solved it.

OpenStudy (anonymous):

I know the answers to these.. I just wanted to know the steps.. so I can actually learn it.

OpenStudy (wolf1728):

Then I should continue with my explanation?

OpenStudy (anonymous):

yes please. :)

OpenStudy (wolf1728):

x^2+5x+6=0 Look for the factors of 6 2 & 3 and 6 & 1 Then I look at the second term 5x and it seems the 2 and the three multiplied equal 6 and 2 + 3 = 5 So, this was an easily factorable equation.

OpenStudy (anonymous):

Thankyou!!!

OpenStudy (wolf1728):

I just thought I'd mention that whenever I factor an equation, I always make sure that the factor of the x² term is always 1. It makes life much simpler. So if you have 3x² + 6x +12 =0 I'd divide it all by 3 x² + 2x + 4 = 0

OpenStudy (anonymous):

Okay- A good tip to remember. thanks.